# Electron-phonon-dominated charge-density-wave fluctuations in TiSe2 accessed by ultrafast nonequilibrium dynamics

### v1

@@ Abstract
1T-TiSe2 exhibits a charge-density-wave (CDW) transition below 200 K, which is believed to be driven by a hybrid exciton-phonon mechanism, making it a versatile platform for investigating the interplay between electronic and lattice degrees of freedom. Although the corresponding band structure modifications below the CDW transition temperature are well established, only a few reports discuss the occurrence of the CDW fluctuating phase and its spectral features well above the transition temperature. Here, we report the direct observation of spectral features associated with CDW fluctuations at 295 K, using time-resolved extreme ultraviolet momentum microscopy. We investigated the transient melting and recovery of CDW fluctuations upon nonresonant ultrafast photoexcitation. Surprisingly, our results reveal that the coherent amplitude mode modulating the ultrafast CDW recovery persists at these elevated temperatures. The time-, energy- and momentum-resolved photoemission data supported by density functional perturbation theory further confirm that CDW fluctuations at these elevated temperatures are dominated by electron-phonon interaction. The analysis of these very localized microscopic fluctuations consequently provides new insights into the complex interplay between the electronic and lattice degrees of freedom at elevated temperatures and, therefore, on the nature of this quantum phase transition.
I Introduction
@@
Understanding the interplay between electronic and lattice degrees of freedom is essential for comprehending phase transitions, transport properties, and the underlying mechanism of complex quantum phases. Disentangling the origin of the charge density wave (CDW) phase in 1T-TiSe2 provides the right platform to do so. Besides the fundamental importance, correct interpretation of this mechanism could unleash the full potential of unconventional CDW-related features, such as CDW chirality Ishioka et al. (2010), tunable nonlinear optical response Zhang et al. (2022), optical gyrotropy Xu et al. (2020), superconductivity Morosan et al. (2006), and an incoherent fluctuating phase Holt et al. (2001). 1T-TiSe2 is a transition metal dichalcogenide (TMDC) that transforms into a CDW phase below K, accompanied by a 2×2×2 lattice reconstruction Di Salvo et al. (1976); Woo et al. (1976a). Despite extensive experimental and theoretical investigations, the driving mechanism of the CDW phase remains a debated question Rossnagel (2011). One perspective attributes the origin to electron-hole coupling, as in the excitonic insulator scenario Di Salvo et al. (1976); Kidd et al. (2002); Cercellier et al. (2007); Monney et al. (2009, 2010, 2012); Kogar et al. (2017), explained by its semiconducting nature and small indirect band gap Watson et al. (2019), which can lead to the spontaneous formation of excitons and a corresponding new ground state Jérome et al. (1967). The alternative perspective argues that the dominant mechanism is electron-phonon coupling, represented by the band-type Jahn-Teller effect Hughes (1977); Yoshida and Motizuki (1980); Whangbo and Canadell (1992); Rossnagel et al. (2002); Weber et al. (2011). Recently, a series of ultrafast time-resolved studies made an effort to disentangle the two contributions Porer et al. (2014); Monney et al. (2016); Karam et al. (2018); Hedayat et al. (2019); Lian et al. (2020); Otto et al. (2021); Cheng et al. (2022); Heinrich et al. (2023), and reach a consensus that electron-hole and electron-lattice interactions are intertwined in driving the CDW transition van Wezel et al. (2010). The ultrafast femtosecond dynamics (i.e., melting and recovery) of the order parameter (e.g., the CDW gap) Rohwer et al. (2011); Hellmann et al. (2012) was used as an argument for the former, while the laser-induced coherent phonon motion of the CDW amplitude mode Porer et al. (2014); Monney et al. (2016); Hedayat et al. (2019); Duan et al. (2021) for the latter interaction.
@@
Several studies have reported that the CDW phase persists above due to the fluctuation of the order parameter related to enigmatic CDW fluctuations Woo et al. (1976b); Miyahara et al. (1995); Holt et al. (2001); Kidd et al. (2002); Chen et al. (2016); Cercellier et al. (2007); Monney et al. (2010, 2016), where the presence of finite correlation lengths leads to phase incoherent CDW domains across the sample. Despite recent reports discussing the CDW fluctuations in 1T-TiSe2 at such elevated temperatures Mizukoshi et al. (2023); Zhang et al. (2022); Cheng et al. (2022); Guo et al. (2025) with some reporting CDW fluctuation signatures in static angle-resolved photoemission spectroscopy (ARPES) Jaouen et al. (2019), none of these directly mapped the nonequilibrium light-induced dynamics of these fluctuations in energy and momentum space. Applying the latter technique to explore this exotic phase could provide fertile ground to scrutinize the interplay and robustness of electron-hole and electron-phonon interactions at elevated temperatures. Indeed, nonequilibrium photoinduced dynamics using time-resolved ARPES (tr-ARPES) have significantly contributed to the ongoing debate on the microscopic origin of CDW fluctuations. Similarly to the observation in the CDW regime, ultrafast sub-100 fs melting of the fluctuating phase, marked by the disappearance of the faint backfolded Se 4p band, was used as evidence for the exciton-driven CDW fluctuations Monney et al. (2016). Contrary to these claims, a recent time-dependent density functional theory study showed that exciton order is melted above Lian et al. (2019). Moreover, the coherent oscillation of the CDW amplitude mode was observed under ambient conditions using time-resolved optical spectroscopy Mizukoshi et al. (2023), suggesting a lattice involvement in the dynamics of CDW fluctuations. As evidenced by these opposing claims, the CDW fluctuating phase in 1T-TiSe2 is still an elusive puzzle, while it would be highly valuable to understand its origin considering its recently discovered connection to unconvetional superconductivity Joe et al. (2014); Jaouen et al. (2019); Hinlopen et al. (2024), as well as its resemblance to the fluctuations preceding ordered phases (e.g., pseudogap) in other TMDCs Borisenko et al. (2008) and strongly correlated materials, such as cuprates Arpaia et al. (2019) and kagome metals Zhong et al. (2024); Liu et al. (2025).
@@
In this work, we use time-resolved extreme ultraviolet (XUV) momentum microscopy to investigate CDW fluctuations in 1T-TiSe2 and their ultrafast light-induced dynamics directly in energy-momentum space. The band structure mapping reveals that the backfolded Se 4p band associated with the CDW phase remains detectable even at room temperature (RT), in agreement with previous reports Kidd et al. (2002); Cercellier et al. (2007); Jaouen et al. (2019). We also measured the partial melting of these CDW fluctuations on an ultrafast timescale, followed by a rapid recovery ( fs). This ultrafast melting signal of CDW fluctuations is offset from the maximum of the hot electron temperature by around fs, and overlaps with the population dynamics of excited Ti 3d states. Together, this demonstrates that CDW fluctuations are able to withstand a strong laser excitation and are not directly linked to the hot electron temperature, suggesting an involvement of the phonon mechanism. Interestingly, we also observe that the recovery of the CDW fluctuations is modulated by coherent CDW amplitude phonon modes. We challenge the belief that the fluctuating phase is ruled solely by the excitonic correlations Monney et al. (2016) by reporting the periodic modulations of the backfolded band spectral weight at RT together with density-functional-theory (DFT) based calculations, indicating not only the preserved electron-phonon coupling mechanism at elevated temperatures but also confirming this to be the main contributor in the hybrid mechanism. More specifically, our calculations incorporating the dynamical electron-phonon scattering confirm that the strongly-coupled CDW-related soft phonon is relevant for the observed equilibrium and transient spectral features of CDW fluctuations. Namely, strong interaction between the soft (yet stable) phonon and Se 4p states leads to the fluctuating replica, while light-induced excitations harden the phonon and screen this interaction, consequently quenching the electron-phonon-driven CDW fluctuations. More generally, our study corrects and directs the emerging consensus on the hybrid exciton-phonon-governed mechanism, with possible implications for the regular CDW phase.
II Equilibrium and Nonequilibrium Electronic Band Structure Mapping
@@
The experiments were performed at the Centre Lasers Intenses et Applications (CELIA) in Bordeaux, France. The time-resolved momentum microscopy setup featuring a polarization-tunable infrared (1.2 eV) pump and XUV probe (21.6 eV) pulses [Fig. 1(a)] allows for mapping of the occupied and unoccupied states (up to 1.2 eV above the Fermi level) over the full photoemission horizon [Fig. 1(b)]. More details about the experimental setup can be found elsewhere Fragkos et al. (2025); Tkach and Schönhense (2025); Tkach et al. (2024a) and in the Methods section.
@@
To gain further insight into the signatures of RT CDW fluctuations, we perform an in-depth analysis of the spectral features around the M(L) pocket for unexcited [Figs. 2 (a)-(b)] and photoexcited 1T-TiSe2 (at pump-probe temporal overlap) [Figs. 2 (c)-(d)]. The photoemission intensity along the (A) - M(L) cut [Fig. 2(a)] reveal spectral features associated with RT CDW fluctuations. First, we note the presence of the well-established backfolded Se 4p valence band, typically used as the ARPES evidence of the emergence of CDW order, exhibited by a faint yet discernible down-dispersive band feature at the M(L) point below the Ti 3d band in Fig. 2(a). This feature is more apparent in the momentum distribution curves (MDC) around Å-1, as indicated by the black bars in Fig. 2(b). The distance between the two peaks increases with higher binding energies, demonstrating the hole-like Se 4p band character. The backfolded Se 4p feature is also observed at pump-probe overlap in both the photoemission intensity image and MDC cut [Figs. 2(c) and (d)], indicating that CDW fluctuations are not completely melted and persist upon photoexcitation at this relatively high fluence (0.95 mJ/cm2). Additionally, a first-principles theoretical study of the electronic response demonstrated that the excitonic mode is quenched upon heating above Lian et al. (2019), while ultrafast pump-probe optical spectroscopy and tr-ARPES studies of 1T-TiSe2 in the CDW regime () argued that the excitonic correlations should be destroyed above the critical fluence of 0.04 mJ/cm2 Porer et al. (2014); Huber et al. (2022); Cheng et al. (2022). This robustness of the fluctuating replica band at RT and high fluence conditions suggests that the CDW fluctuations are not solely of excitonic origin.
III Ultrafast Melting and Recovery of RT-CDW Fluctuations
@@
Now that we have established the equilibrium spectral signatures associated with RT-CDW fluctuations, we turn our attention to their ultrafast melting and recovery upon photoexcitation to gain detailed insight into the microscopic origin underlying the formation of this exotic phase Boschini et al. (2024). The tr-ARPES enables systematic temporal separation of fundamental electronic and structural processes, allowing the identification of the dominant interactions Hellmann et al. (2012).
@@
In this spirit, we record the three-dimensional photoemission intensity as a function of pump-probe delay (between -0.5 ps to 3 ps), yielding a four-dimensional dataset . From this multidimensional dataset, one can hunt for subtle signatures of RT CDW fluctuations melting and recovery along energy, momentum, and time axes. We start by looking at the differential (pre-time-zero signal subtracted) CEC near the Fermi level ( = -0.07 eV), at a time delay of [Figs. 3 (a) and (b)]. These CECs reveal the population of the conduction Ti 3d band (hollow elliptical positive differential signal) and suppression of the CDW backfolded band spectral weight (central negative differential signal) upon photoexcitation, around M(L) pockets. At this binding energy, after 400 fs, the negative differential signal associated with the melting of CDW order vanishes, as the filling of the bottom of the conduction band becomes dominant. This behavior is also evidenced in the differential energy-momentum region around the L(M) point, extracted for the same pump-probe delays.
@@
To get more insight into the concomitant dynamics of hot electrons and CDW melting/recovery, we analyze the time-resolved photoemission intensities [Fig. 3(c)] in selected regions-of-interest (ROIs), shown as colored boxes in Fig. 3(b). The purple dots represent the IR/XUV cross-correlation extracted from the time-dependent laser-assisted photoemission signal (ROI not shown), with an FWHM duration of 171 fs (Gaussian fit, yellow line). The dark red curve represents hot electron populations in Ti 3d conduction band, showing a fast light-induced rise followed by a single exponential decay behavior. The timescale of this relaxation process is strongly dependent on the excess energy of these hot electrons, as shown in the Fig. S1 in Supplemental Material (SM). As discussed in the context of the analysis of the CECs, the black ROI is spectrally congested, with some overlapping contribution of both the bottom of the Ti 3d conduction band and the top of the Se 4p CDW backfolded band. The dynamics in this ROI feature a fast decay (experimental-resolution-limited) associated with the melting of CDW fluctuations, followed by a refilling and subsequent decay associated with the intraband hot electron dynamics.
@@
Another interesting feature in the time-resolved photoemission signal is evident in panel 3(e), where we compare the dynamics of the backfolded Se 4p band with the time evolution of the electronic temperature (extracted from the Fermi-Dirac distribution fitting) and of the total excited state population taken from the yellow ROI in Fig. 3(b), representing the excitation and thermalization of Ti 3d states in a wider energy and momentum region. Namely, the minimum of the backfolded band intensity is offset from the maximum of the electronic temperature, while it approximately follows the early dynamics of the excited state population of Ti 3d bands. This comparison suggests a mechanism beyond the excitonic insulator scenario, considering that the excitonic order is expected to be melted when the electron temperature reaches its maximum. The concomitant observation of partial melting under relatively high-fluence condition (0.95 mJ/cm2), the finite delay between the peak of transient electronic temperature and quenching of the backfolded Se 4p band, as well as coherent CDW amplitude phonon mode modulating the recovery dynamics, supports the idea of the dominant phonon contribution in the assumed hybrid exciton and electron-phonon mechanism at such elevated temperatures. The idea of dominant phonon contribution is further supported by a recent study Buchberger et al. (2025) which shows that while the band gap is sensitive to substrate screening, the temperature-dependent evolution of its electronic structure remains unaffected, indicating that excitons are not a primary mechanism that drives the CDW transition in TiSe2.

### v3

@@ Abstract
The complex phase diagram of 1T-TiSe2 consists of a charge density wave (CDW) below 200 K, and CDW fluctuations of still unknown origin at higher temperatures. Here, we use time-resolved extreme ultraviolet momentum microscopy and density functional perturbation theory to uncover the formation mechanism of CDW fluctuations and their spectral features at 295 K. We investigated the transient dynamics of fluctuations upon nonresonant ultrafast photoexcitation, and directly correlate it with the CDW soft-phonon hardening. Surprisingly, our results show that the coherent amplitude mode modulating ultrafast CDW recovery persists above , and reveal that CDW fluctuations are dominated by the electron-phonon interaction rather than excitonic correlations as commonly believed. Our findings on these microscopic CDW fluctuations clarify the complex interplay between electronic and lattice degrees of freedom at elevated temperatures and, therefore, could be useful in understanding the nature of the CDW phase transition in 1T-TiSe2 and similar quantum materials.
Introduction
@@
Understanding the interplay between electronic and lattice degrees of freedom is essential for comprehending phase transitions, superconductivity, transport properties, and the underlying mechanism of complex quantum phases. Disentangling the origin of the charge density wave (CDW) phase in 1T-TiSe2, with K and a lattice reconstruction Di Salvo et al. (1976); Woo et al. (1976a), provides the right platform to do so. Despite extensive experimental and theoretical investigations, the driving mechanism of the CDW phase remains a debated question Rossnagel (2011). One perspective attributes the origin to electron-hole coupling Jérome et al. (1967), as in the excitonic insulator scenario Di Salvo et al. (1976); Kidd et al. (2002); Cercellier et al. (2007); Monney et al. (2009, 2010, 2012); Kogar et al. (2017), while alternative argues that the dominant mechanism is electron-phonon coupling, represented by the band-type Jahn-Teller effect Hughes (1977); Yoshida and Motizuki (1980); Whangbo and Canadell (1992); Rossnagel et al. (2002); Weber et al. (2011). Recently, a series of ultrafast time-resolved studies made an effort to disentangle the two contributions Porer et al. (2014); Monney et al. (2016); Karam et al. (2018); Hedayat et al. (2019); Lian et al. (2020); Otto et al. (2021); Cheng et al. (2022); Heinrich et al. (2023), and reach a consensus that electron-hole and electron-lattice interactions are intertwined in driving the CDW transition van Wezel et al. (2010). The ultrafast femtosecond dynamics (i.e., melting and recovery) of the order parameter (e.g., the CDW gap) Rohwer et al. (2011); Hellmann et al. (2012) was used as an argument for the former, while the laser-induced coherent phonon motion of the CDW amplitude mode Porer et al. (2014); Monney et al. (2016); Hedayat et al. (2019); Duan et al. (2021) for the latter interaction.
@@
Several studies have reported that the CDW phase persists above due to the fluctuation of the order parameter related to enigmatic CDW fluctuations Woo et al. (1976b); Miyahara et al. (1995); Holt et al. (2001); Kidd et al. (2002); Chen et al. (2016); Cercellier et al. (2007); Monney et al. (2010, 2016), where the presence of finite correlation lengths leads to phase incoherent CDW domains across the sample. Despite recent reports discussing the CDW fluctuations in 1T-TiSe2 at such elevated temperatures Mizukoshi et al. (2023); Zhang et al. (2022); Cheng et al. (2022); Guo et al. (2025) with some reporting CDW fluctuation signatures in static angle-resolved photoemission spectroscopy (ARPES) Jaouen et al. (2019), none of these directly mapped the nonequilibrium light-induced dynamics of these fluctuations in energy and momentum space. Applying the latter technique to explore this exotic phase could provide fertile ground to scrutinize the interplay and robustness of electron-hole and electron-phonon interactions at elevated temperatures. Indeed, a study using time-resolved ARPES (tr-ARPES) demonstrated sub-100 fs melting of the fluctuating phase, which was used as an evidence for the exciton-driven CDW fluctuations Monney et al. (2016). Contrary to these claims, a recent time-dependent density functional theory study showed that exciton order is melted above Lian et al. (2019). Moreover, the coherent oscillation of the CDW amplitude mode was observed under ambient conditions using time-resolved optical spectroscopy Mizukoshi et al. (2023), suggesting a lattice involvement in the dynamics of CDW fluctuations. As evidenced by these opposing claims, the CDW fluctuating phase in 1T-TiSe2 is still an elusive puzzle, while it would be highly valuable to understand its origin considering its recently discovered connection to unconvetional superconductivity Joe et al. (2014); Jaouen et al. (2019); Hinlopen et al. (2024), as well as its resemblance to the fluctuations preceding ordered phases (e.g., pseudogap) in other transition metal dichalcogenides (TMDCs) Borisenko et al. (2008) and strongly correlated materials, such as cuprates Arpaia et al. (2019) and kagome metals Zhong et al. (2024); Liu et al. (2025). Unraveling nonequilibrium properties of CDW fluctuations in TiSe2 would also be highly useful for comprehending recently discovered light-induced hidden phases in TiSe2 Duan et al. (2023); Huber et al. (2024).
@@
In this work, we resolve the debate on the origin of CDW fluctuations in 1T-TiSe2 by using time-resolved extreme ultraviolet (XUV) momentum microscopy, which allows ultrafast light-induced dynamics directly in energy-momentum space. The band structure mapping reveals that the backfolded Se 4p band associated with the CDW phase remains detectable even at room temperature (RT), in agreement with previous reports Kidd et al. (2002); Cercellier et al. (2007); Jaouen et al. (2019). We also measured the partial melting of these CDW fluctuations on an ultrafast timescale, followed by a rapid recovery ( fs). This ultrafast melting signal of CDW fluctuations is offset from the maximum of the hot electron temperature by around fs, and overlaps with the population dynamics of excited Ti 3d states. Together, this demonstrates that CDW fluctuations are able to withstand a strong laser excitation and are not directly linked to the hot electron temperature, suggesting an involvement of the phonon mechanism. Interestingly, we also observe that the recovery of the CDW fluctuations is modulated by coherent CDW amplitude phonon modes. We challenge the belief that the fluctuating phase is ruled solely by the excitonic correlations Monney et al. (2016) by reporting the periodic modulations of the backfolded band spectral weight at RT together with density-functional-theory (DFT) based calculations, indicating not only the preserved electron-phonon coupling mechanism at elevated temperatures but also confirming this to be the main contributor in the hybrid mechanism. More specifically, our calculations incorporating the dynamical electron-phonon scattering confirm that the strongly-coupled CDW-related soft phonon is relevant for the observed equilibrium and transient spectral features of CDW fluctuations. Namely, strong interaction between the soft (yet stable) phonon and Se 4p states leads to the fluctuating replica, while light-induced excitations harden the phonon and screen this interaction, consequently quenching the electron-phonon-driven CDW fluctuations. More generally, our study poses a series of strong arguments favoring the electron-phonon mechanism of CDW fluctuation formation at elevated temperature, which also puts a strong emphasis on the electron-phonon scenario in the debate on the origin of the regular CDW phase. Besides the fundamental importance, correct interpretation of this mechanism could unleash the full potential of unconventional CDW-related features, such as CDW chirality Ishioka et al. (2010), tunable nonlinear optical response Zhang et al. (2022), optical gyrotropy Xu et al. (2020), and superconductivity Morosan et al. (2006)
Results
@@
Equilibrium and Nonequilibrium Electronic Band Structure Mapping. The experiments were performed at the Centre Lasers Intenses et Applications (CELIA) in Bordeaux, France. The time-resolved momentum microscopy setup featuring a polarization-tunable infrared (1.2 eV, sub-150) pump and XUV probe (21.6 eV, sub-50) pulses [Fig. 1(a)] allows for mapping of the occupied and unoccupied states (up to 1.2 eV above the Fermi level) over the full photoemission horizon [Fig. 1(b)]. More details about the experimental setup can be found elsewhere Fragkos et al. (2025); Tkach and Schönhense (2025); Tkach et al. (2024a) and in the Methods section.
@@
To gain further insight into the signatures of RT-CDW fluctuations, we perform an in-depth analysis of the spectral features around the M(L) pocket for unexcited [Figs. 2 (a)-(b)] and photoexcited 1T-TiSe2 (at pump-probe temporal overlap) [Figs. 2 (c)-(d)]. The photoemission intensity along the (A) - M(L) cut [Fig. 2(a)] reveal spectral features associated with RT-CDW fluctuations. First, we note the presence of the well-established backfolded Se 4p valence band, typically used as the ARPES evidence of the emergence of CDW order, exhibited by a faint yet discernible down-dispersive band feature at the M(L) point below the Ti 3d band in Fig. 2(a). This feature is more apparent in the momentum distribution curves (MDC) around Å-1, as indicated by the black bars in Fig. 2(b). The distance between the two peaks increases with higher binding energies, demonstrating the hole-like Se 4p band character. The backfolded Se 4p feature is also observed at pump-probe overlap in both the photoemission intensity image and MDC cut [Figs. 2(c) and (d)], indicating that CDW fluctuations are not completely melted and persist upon photoexcitation at this relatively high fluence (0.95 mJ/cm2). Additionally, a first-principles theoretical study of the electronic response demonstrated that the excitonic mode is quenched upon heating above Lian et al. (2019), while ultrafast pump-probe optical spectroscopy and tr-ARPES studies of 1T-TiSe2 in the CDW regime () argued that the excitonic correlations should be destroyed above the critical fluence of 0.04 mJ/cm2 Porer et al. (2014); Huber et al. (2022); Cheng et al. (2022). This robustness of the fluctuating replica band at RT and high fluence conditions suggests that the CDW fluctuations are not solely of excitonic origin.
@@
Ultrafast Melting and Recovery of RT-CDW Fluctuations. Now that we have established the equilibrium spectral signatures associated with RT-CDW fluctuations, we turn our attention to their ultrafast melting and recovery upon photoexcitation to gain detailed insight into the microscopic origin underlying the formation of this exotic phase Boschini et al. (2024). The tr-ARPES enables systematic temporal separation of fundamental electronic and structural processes, allowing the identification of the dominant interactions Hellmann et al. (2012).
@@
In this spirit, we record the three-dimensional photoemission intensity as a function of pump-probe delay (between -0.5 ps to 3 ps), yielding a four-dimensional dataset . From this multidimensional dataset, one can hunt for subtle signatures of RT-CDW fluctuations melting and recovery along energy, momentum, and time axes. We start by looking at the differential (pre-time-zero signal subtracted) CEC near the Fermi level ( = -0.07 eV), at a time delay of [Figs. 3 (a) and (b)]. These CECs reveal the population of the conduction Ti 3d band (hollow elliptical positive differential signal) and suppression of the CDW backfolded band spectral weight (central negative differential signal) upon photoexcitation, around M(L) pockets. At this binding energy, after 400 fs, the negative differential signal associated with the melting of CDW order vanishes, as the filling of the bottom of the conduction band becomes dominant. This behavior is also evidenced in the differential energy-momentum region around the L(M) point, extracted for the same pump-probe delays.
@@
To get more insight into the concomitant dynamics of hot electrons and CDW melting/recovery, we analyze the time-resolved photoemission intensities [Fig. 3(c)] in selected regions-of-interest (ROIs), shown as colored boxes in Fig. 3(b). The purple dots represent the IR/XUV cross-correlation extracted from the time-dependent laser-assisted photoemission signal (ROI not shown, cross-correlation shown in Supplementary Figure 6), with an FWHM duration of 171 fs (Gaussian fit, yellow line). The dark red curve represents hot electron populations in Ti 3d conduction band, showing a fast light-induced rise followed by a single exponential decay behavior. The timescale of this relaxation process is strongly dependent on the excess energy of these hot electrons, as shown in the Supplementary Figure 1 in Supplementary Information (SI). As discussed in the context of the analysis of the CECs, the black ROI is spectrally congested, with some overlapping contribution of both the bottom of the Ti 3d conduction band and the top of the Se 4p CDW backfolded band. The dynamics in this ROI feature a fast decay (experimental-resolution-limited) associated with the melting of CDW fluctuations, followed by a refilling and subsequent decay associated with the intraband hot electron dynamics.
@@
Another interesting feature in the time-resolved photoemission signal is evident in panel 3(e), where we compare the dynamics of the backfolded Se 4p band with the time evolution of the electronic temperature (extracted from the Fermi-Dirac distribution fitting) and of the total excited state population taken from the yellow ROI in Fig. 3(b), representing the excitation and thermalization of Ti 3d states in a wider energy and momentum region. Note that the increase in electronic temperature and excited-state population observed at negative time delays is attributed to a weak prepulse extending up to a few hundred femtoseconds in the temporal profile of the infrared pump pulses. More importantly, the minimum of the backfolded band intensity is offset from the maximum of the electronic temperature, while it approximately follows the early dynamics of the excited state population of Ti 3d bands. This comparison suggests a mechanism beyond the excitonic insulator scenario, considering that the excitonic order is expected to be melted when the electron temperature reaches its maximum. Namely, exciton, if it exists, would be formed from the holes in Se 4p and electrons in Ti 3d bands, and, therefore, its quenching should be correlated with the total photo-excited carrier density, which can be described with the effective temperature. Furthermore, it is expected that the suppression of the excitonic mode is very fast and that it should correspond to the fastest signal, however, the RT-CDW signal suppression is offset from by approximately fs.
The concomitant observation of partial melting under relatively high-fluence condition (0.95 mJ/cm2), the finite delay between the peak of transient electronic temperature and quenching of the backfolded Se 4p band, as well as coherent CDW amplitude phonon mode modulating the recovery dynamics, supports the idea of the dominant phonon contribution in the assumed hybrid exciton and electron-phonon mechanisms at such elevated temperatures. The idea of dominant phonon contribution is further supported by a recent study Buchberger et al. (2025) which shows that while the band gap is sensitive to substrate screening, the temperature-dependent evolution of its electronic structure remains unaffected, indicating that excitons are not a primary mechanism that drives the CDW transition in TiSe2.
Discussion